OpenStudy (anonymous):

1/2= a/9 The directions are- Find the value of the letter in each proportion

OpenStudy (anonymous):

1/2= a/9 a = 9*(1/2) a = 9/2 a = 4.5

OpenStudy (anonymous):

you have to isolate a by deviding bothe sides by 1/9. so 1/9/1/9a=1/2/1/9 deviding bothe sids yeilds a=9/2

OpenStudy (anonymous):

1/2=a/9 a=1/2*9 =9/2=4.5